M 408D 56485,90,95 Spring ’06 TTh 12:30-2 WRW 102; MW 9-10 CPE 2.206 (56485), MW 1-2 RLM 7.120 (56490), MW 3-4 RLM 6.120 (56495)

*** T Th classes will be lectures, M W discussion sections.

Instructor: C. Friedman RLM 11.140 471-5161 office hrs: T Th 2:30-3:30pm or by appt. Web Page (homework assignments are listed): www.ma.utexas.edu/˜ friedman

Text and Syllabus: Calculus, Stewart 5th ed. Syllabus: 7.7, 8.8, 12.1-12.10, 12.12, 11.1-11.4, 13.1-13.6, 14.1-14.2, 15.1-15.8, 16.1-16.7 (Subject to change.)

Exams: We will have exams on 2/21 (class 11; Tues.), 4/4 (class 21; Tues.), 5/4 (Class 30; Thurs.) (No “Final exam” planned [subject to Department approval]). All exam grades will be curved. MAKEUPS WILL BE GIVEN VERY RELUCTANTLY

Homework: Homework will be assigned each Tuesday in class and will be due the next Monday in Discussion Section. A serious effort should be made on these exercises. It will probably be difficult to completely learn the course material without working problems; also exam questions may bear some resemblance to homework exercises. Additionally, all homework together will be counted as a grade equivalent to one exam grade.

Quizes: Three quizes will be given in the discussion classes on 2/13 (Mon.), 3/27 (Mon.), and 4/26 (Wed.). These will probably be short (about 1/2 hour) and consist of one problem. All three quiz grades will be combined to form one grade counted equal to a single test grade.

Final Grades: In computing a final grade, I will count the three tests, one total homework grade and one total quiz grade equally. (So there will be five grades to average.)